# Large Language Models with Task-Oriented Experts

### v1

@@ Abstract
Construction of a general-purpose post-recognition error corrector poses a crucial question: how can we most effectively train a model on a large mixture of domain datasets? The answer would lie in learning dataset-specific features and digesting their knowledge in a single model. Previous methods achieve this by having separate correction language models, resulting in a significant increase in parameters. In this work, we present Mixture-of-Experts as a solution, highlighting that MoEs are much more than a scalability tool. We propose a Multi-Task Correction MoE, where we train the experts to become an “expert” of speech-to-text, language-to-text and vision-to-text datasets by learning to route each dataset’s tokens to its mapped expert. Experiments on the Open ASR Leaderboard show that we explore a new state-of-the-art performance by achieving an average relative $5.0$% WER reduction and substantial improvements in BLEU scores for speech and translation tasks. On zero-shot evaluation, NeKo outperforms GPT-3.5 and Claude-Opus with $15.5$% to $27.6$% relative WER reduction in the Hyporadise benchmark. NeKo performs competitively on grammar and post-OCR correction as a multi-task model.
[Uncaptioned image]NeKo: Toward Post Recognition Generative Correction
Large Language Models with Task-Oriented Experts
Yen-Ting Linthanks: Work done at NVIDIA research as an intern. Chao-Han Huck Yang  Zhehuai Chen Piotr Zelasko Xuesong Yang Zih-Ching Chen Krishna C Puvvada Szu-Wei Fu  Ke Hu  Jun Wei Chiu  Jagadeesh Balam  Boris Ginsburg  Yu-Chiang Frank Wang NVIDIA corresponding authors: ytl@ieee.org, hucky@nvidia.com
@@ 1 Introduction
Human recognition (Biederman, 1987; Juang and Furui, 2000; Kanwisher et al., 1996) capabilities span multiple modalities, including speech recognition, visual patterns, and extensions to semantic and textual interpretations. These faculties, however, are not infallible and often incorporate mis-recognition errors. Despite these imperfections, humans efficiently communicate using speech, language, or facial expressions.
@@
For instance, two non-native speakers (Lev-Ari, 2015; Valaki et al., 2004) can often achieve mutual understanding through this imperfect recognition and subsequent interpretative processes, even when the conversation is marred by lexical inaccuracies and subdued accents. In other words, humans (as intelligent agents) exhibit a robust capacity for generative understanding Jiang et al. (2020); Cheng et al. (2021) that extends beyond initial recognition results.
In neuroscience Zatorre and Gandour (2008), the inferior temporal gyrus and the temporal lobe are not confined to rudimentary perception but are also integral to the post-recognition processes that facilitate semantic understanding of language (Levinson and Evans, 2010), speech (Marshall et al., 2015), and visual patterns (Vink et al., 2020). This form of “post-recognition correction,” exemplified by the application of language modeling (LM) to initial recognition outputs, has been introduced to the field for both acoustic (automatic speech recognition, ASR) and visual (optical character recognition, OCR) modalities since the early explorations (Jelinek, 1976; Dixon and Silverman, 1975) of learning algorithms in 1970s. The most prevalent approaches to utilizing LMs for post-recognition boosting are predominantly ranking or retrieval-based. In these setups, the LM is tasked with ranking and scoring (Ljolje et al., 1999) the top $n$-best hypotheses generated by the first-pass recognition system. This process often incorporates “discriminative modeling” algorithms (Sukkar and Lee, 1996; Mangu et al., 2000) and representation embeddings, such as BERT (Salazar et al., 2020; Kenton and Toutanova, 2019), to minimize the word error rate (WER) during training (Prabhavalkar et al., 2018; Mangu et al., 2000).
@@
With the LMs scaling up to LLMs (Brown et al., 2020), recent efforts (Chan et al., 2023; Yang et al., 2023; CHEN et al., 2023; Hu et al., 2024a) have focused on exploring a “generative modeling” for post-recognition correction. This generative error correction (GER) approach uses LLMs to conduct final recognition from given first-pass text-based predictions from recognition models, including ASR, image captioning (IC), and machine translation (MT). This cascaded two-agents text-to-text GER model has outperformed larger single multi-modal and multi-task models in these tasks.
Meanwhile, these GER solutions heavily depends on domain-specific fine-tuning processes (Chen et al., 2024a) that utilize parameter-efficient components, which often suffers a performance degradation from a lack of generalizability across different datasets, domains, and tasks. In other words, how to design and further openly provide a “general-purpose post (every) recognition correction model” is still one undiscovered and crucial topic within the research community.
On the other hand, directly fine-tuning LLMs on a mixture of diverse error correction datasets can lead to suboptimal performance (CHEN et al., 2023; Lange et al., 2022) due to differences in input modalities, output formats, error types, and domain characteristics. For example, ASR errors stem from phonetic similarities or acoustic ambiguities, while OCR errors involve visual or character-level confusions. Additionally, error distributions can vary widely across datasets, even within the same task.
@@
To characterize “model generalization,” mixture-of-experts (MoE) (Jiang et al., 2024a) has emerged as a promising approach for multi-task learning, consisting of of a set of expert networks and a gating network that learns to route the input to the most appropriate expert(Sukhbaatar et al., 2024). This enables MoE models to learn more specialized and fine-grained representations compared to monolithic models. However, most MoE models are designed for general-purpose language modeling(Dai et al., 2024), with experts not explicitly assigned to specific tasks, but rather learn to specialize in different aspects of the input space through data-driven training, expect for a recent vision work (Ye and Xu, 2023). Effectively leverage MoE for multi-task error correction, where the experts need to capture task-specific features while allowing knowledge sharing, remains an open question.
@@
In this work, we propose NeKo, a “geNerative multi-tasK error correction” approach that leverages a pre-trained MoE model to drive diverse tasks and cross-domain knowledge, as shown in Figure 1. The key idea is to continuously pre-train MoE model on a mixture of error correction datasets, with each expert specializing in a specific domain. This task-oriented MoE fine-tuning approach enables the experts to capture task-specific features while allowing knowledge sharing through the router (Dai et al., 2024).
@@
NeKo captures the nuances of each task, benefiting from shared knowledge across experts. Evaluated on tasks such as ASR, ST, OCR, and unseen textual error correction (TEC), NeKo consistently outperforms baseline models, including Claude-Opus and GPT-3.5. It achieves state-of-the-art WER reduction on the Hyporadise benchmark and large-scale Open ASR Leaderboard (Srivastav et al., 2023). NeKo also significant improves in OCR error correction. Further analysis confirms its robust multi-task capabilities.
In summary, the main contributions of this work include:
@@
    We introduce NeKo, a multi-task error correction LLM that leverages task-oriented mixture-of-experts for diverse post-recognition correction tasks. To the best of our knowledge, this is the first work that explores the use of MoE for multi-task error correction.
@@
    NeKo has been studied under a new form of cross-modalities post-recognition correction evaluation, serving as strong open-source ASR, ST, OCR, and TEC baselines. Our results show that NeKo discovers new state-of-the-art performance in ASR as a multi-task error correction model.
@@
    The NeKo models, newly created source datasets, and training processes are scheduled to open source under the CC BY-SA 4.0 license to support reproducibility and to encourage future research.
2 Related Work
Language Modeling and Generative Error Correction
Neural correction LMs have been widely used for end-to-end (E2E) models for text error correction or normalization for both ASR (Jelinek, 1990; Irie et al., 2016; Zhang et al., 2019a, b; Guo et al., 2019) and OCR (Plamondon and Srihari, 2000; Sabir et al., 2017; Fogel et al., 2020). These models often use beam search to generate new estimates, and can usually handle text normalization and denormalization or spelling errors. To utilize the abundant textual data to improve ASR, neural LMs are integrated to E2E models using shallow fusion or rescoring over acoustic model’ likelihood confidence (Kannan et al., 2018; Salazar et al., 2020; Yang et al., 2021).
In addition to using only textual hypotheses for correction, deliberation models (Hu et al., 2020; Wang et al., 2022; Hu et al., 2023a) utilizes both audio and text for generating new ASR hypotheses. While the aforementioned methods have drawn much attention in the past, recent advances focus on using pretrained textual LLMs to benefit E2E ASR (Song et al., 2023; Hu et al., 2023c), or similar to deliberation, employing both speech and text for speech understanding by prompting (Gong et al., 2023) and joint speech and language foundation models (SLMs) (Wang et al., 2023). On the other hand, text-to-text based GEC (CHEN et al., 2023; Yang et al., 2023; Radhakrishnan et al., 2023; Chan et al., 2023) have been recently introduced to ASR, ST, image captioning, and video summarization to set up superior performance  [^1] GEC have shown better results compared to a similar scale of (i) multi-modal injection model and (ii) re-ranking based model, where we will leave these studies in the future upon having one multi-task correction LM as NeKo presented here. Please also. These GEC approaches take the text output as an input to large language models (LLMs), enabling them to perform zero-shot typo correction by generatively refining the final recognition results (CHEN et al., 2023; Yang et al., 2023; Radhakrishnan et al., 2023; Chan et al., 2023). In this work, we aim to design a text-to-text model as a first step for providing open-source post-recognition correction to the communities.
Mixture of Experts
Mixture-of-experts (MoE) (Shazeer et al., 2017) is a machine learning concept that employs multiple expert layers, each of which specializes in solving a specific subtask. The experts then work together to solve the entire task at hand. Recently, MoE has been widely applied to large-scale distributed Deep Learning models by using a cross-GPU layer that exchanges hidden features from different GPUs (Lepikhin et al., 2021; Fedus et al., 2022). The MoE approach is differentiated from existing scale-up approaches for DNNs, such as increasing the depth or width of DNNs, in terms of its high cost-efficiency. Specifically, adding more model parameters (experts) in MoE layers does not increase the computational cost per token at inference time. Thus, MoE has been studied for scaling the models to trillion-size parameters in NLP (Fedus et al., 2022). In speech processing, Gaur et al. (2021); Hu et al. (2023b); You et al. (2021) primarily showed benefits of MoE on speech recognition tasks. In this work, we further pursue this direction by modeling MoE on error correction and highlight the effectiveness and robustness of MoEs in learning from a mixture of correction datasets.
@@ 3.1 Mixture-of-Experts (MoE)
3.2 Task-Oriented Expert Assignment
@@
[…]
@@ 3.3 Training Objective
where $x$ is the input sequence (e.g., ASR hypotheses, OCR output), $y$ is the target sequence (e.g., ground-truth transcription, corrected text), and $p(y|x,T_{i})$ is the probability of the target sequence given the input sequence and the task prompt (Figure 3.) By jointly training on multiple error correction datasets with task-oriented expert assignment, NeKo learns to capture task-specific features while allowing for knowledge sharing across tasks through the shared gating network and other model components.
@@ 4.1 Training and Evaluation Datasets
For the speech translation error correction task, we use the subset of the HypoTranslate dataset (Hu et al., 2024b) for training and evaluation. This dataset includes translation results from FLEURS (Conneau et al., 2022), CoVoST-2 (Wang et al., 2020), and MuST-C (Di Gangi et al., 2019), covering a range of languages such as Spanish, French, Italian, Japanese, Portuguese, Chinese, and Persian. As an extra zero-shot textual correction setup, we evaluate NeKo on machine translation (MT) of WMT’20 for Japanese and Chinese (Barrault et al., 2020a). We use the BLEU score (Papineni et al., 2002) as the evaluation metric for ST (with training and test) and MT (zero-shot).
@@
For the optical character recognition (OCR) error correction task, we use the English portion of the Post-OCR Correction dataset (PleIAs, 2023), which contains newspaper texts from Chronicling America. The dataset includes original texts with varying numbers of OCR mistakes and their corresponding corrected versions. To evaluate our model, we take the first 1,000 characters of both the input text with OCR errors and the ground-truth corrected text. We use the WER as the evaluation metric.
@@
For the textual error correction (TEC) task, we use a subset of the CoEdIT dataset (Raheja et al., 2023), which contains a diverse collection of 82K task-specific instructions for text editing. We select two editing tasks from CoEdIT that align with our error correction objectives: grammar correction and coherence improvement. These tasks focus on correcting grammatical errors and improving the overall coherence of the text, making them suitable for evaluating the effectiveness of our model in handling TEC-related editing instructions. We use the word error rate as the evaluation metric.
@@ 4.2 Task-Specific Recognition Systems and Baselines
We compare against state-of-the-art ASR models, Distil-Whisper-V2-Large(Gandhi et al., 2023), Whisper-V2-Large, Whisper-V3-Large(Radford et al., 2022), Canary(NVIDIA, 2024) without applying GEC method. A end-to-end ASR-LLM, SALM (Chen et al., 2024b), is also compared. For all Cascaded ASR+GEC Methods, the task-specific system is the Canary model. This model transcribes the speech data and generate 5-best hypotheses for each utterance using temperature-based sampling (Ackley et al., 1985) with $p=0.3$. This allows us to capture a diverse set of potential transcriptions for each utterance, which can then be fed into our error correction model.
@@
We compare our proposed method against two baselines: (1) the input text without any correction (denoted as Baseline) and (2) a Mixtral 8x7B model fine-tuned only on the respective dataset for each task (denoted as Mistral 8x7B Direct Finetune). This allows us to assess the effectiveness of our task-oriented expert assignment approach in handling OCR and TEC errors, as well as its ability to leverage knowledge from multiple tasks to improve performance on individual tasks compared to direct fine-tuning on a single dataset.
Table 1: Cross-domain ASR correction results in zero-shot and few-shot settings on the Hyporadise benchmark (CHEN et al., 2023). We compare NeKo against GPT-3.5 Turbo and Claude-Opus in 0- and 5-shot settings. The baseline represents the WER of task-specific model Whisper-Large. The oracle results used in CHEN et al. (2023) (N-best and Compositional) provide an upper bound for the correction performance.
Domain Test Set Baseline GPT-3.5 Turbo Claude-Opus 0-shot w/ NeKo Oracle
@@
Specific Scenario WSJ-dev93 9.0 $8.5_{{\color[rgb]{0,.5,.5}-5.6\%}}$ $7.7_{{\color[rgb]{0,.5,.5}-14.4\%}}$ $8.2_{{\color[rgb]{0,.5,.5}-8.9\%}}$ $7.4_{{\color[rgb]{0,.5,.5}-17.8\%}}$ $8.6_{{\color[rgb]{0,.5,.5}-4.4\%}}$ $7.5_{{\color[rgb]{0,.5,.5}-16.7\%}}$ $\textbf{6.8}_{{\color[rgb]{0,.5,.5}-24.4\%}}$ 6.5 5.3
@@
Common Noise CHiME4-bus 18.8 $17.6_{{\color[rgb]{0,.5,.5}-6.4\%}}$ $16.2_{{\color[rgb]{0,.5,.5}-13.8\%}}$ $17.1_{{\color[rgb]{0,.5,.5}-9.0\%}}$ $15.7_{{\color[rgb]{0,.5,.5}-16.5\%}}$ $17.7_{{\color[rgb]{0,.5,.5}-5.9\%}}$ $15.9_{{\color[rgb]{0,.5,.5}-15.4\%}}$ $\textbf{14.5}_{{\color[rgb]{0,.5,.5}-22.9\%}}$ 16.8 10.7
@@
Speaker Accent MCV-af 25.3 $24.9_{{\color[rgb]{0,.5,.5}-1.6\%}}$ $23.6_{{\color[rgb]{0,.5,.5}-6.7\%}}$ $24.4_{{\color[rgb]{0,.5,.5}-3.6\%}}$ $23.0_{{\color[rgb]{0,.5,.5}-9.1\%}}$ $25.0_{{\color[rgb]{0,.5,.5}-1.2\%}}$ $23.3_{{\color[rgb]{0,.5,.5}-7.9\%}}$ $\textbf{21.0}_{{\color[rgb]{0,.5,.5}-17.0\%}}$ 23.6 21.7
@@
MCV-in 28.6 $27.6_{{\color[rgb]{0,.5,.5}-3.5\%}}$ $25.0_{{\color[rgb]{0,.5,.5}-12.6\%}}$ $27.0_{{\color[rgb]{0,.5,.5}-5.6\%}}$ $24.3_{{\color[rgb]{0,.5,.5}-15.0\%}}$ $27.8_{{\color[rgb]{0,.5,.5}-2.8\%}}$ $24.6_{{\color[rgb]{0,.5,.5}-14.0\%}}$ $\textbf{22.2}_{{\color[rgb]{0,.5,.5}-22.4\%}}$ 27.1 22.6
MCV-sg 26.4 $26.5_{{\color[rgb]{.5,.5,.5}+0.4\%}}$ $25.1_{{\color[rgb]{0,.5,.5}-4.9\%}}$ $25.9_{{\color[rgb]{0,.5,.5}-1.9\%}}$ $24.5_{{\color[rgb]{0,.5,.5}-7.2\%}}$ $26.6_{{\color[rgb]{.5,.5,.5}+0.8\%}}$ $24.7_{{\color[rgb]{0,.5,.5}-6.4\%}}$ $\textbf{22.3}_{{\color[rgb]{0,.5,.5}-15.5\%}}$ 25.5 22.2
Table 2: ASR correction results on the Open ASR Leaderboard. We report the Word Error Rate (WER) for each dataset and the average across all 9 datasets. NeKo establishes a new state-of-the-art performance on the leaderboard, outperforming both end-to-end ASR methods and cascaded ASR+GEC approaches. Lower WER indicates better performance.
Model AMI Earnings22 Gigaspeech LS Clean LS Other SPGI Tedlium Voxp. MCV9 Average
End-to-end ASR Methods
Distil-Whisper-V2-Large (Gandhi et al., 2023) 14.65 12.12 10.31 2.95 6.39 3.28 4.30 8.22 12.60 8.31
Whisper-V2-Large (Radford et al., 2022) 16.82 12.02 10.57 2.56 5.16 3.77 4.01 7.50 10.11 8.06
Whisper-V3-Large 16.01 11.30 10.02 2.03 3.91 2.95 3.90 9.52 9.67 7.70
Canary (NVIDIA, 2024) 14.00 12.25 10.19 1.49 2.49 2.06 3.58 5.81 7.75 6.67
Bestow Speech LM (Chen et al., 2024c) 12.58 12.86 10.06 1.64 3.07 2.11 3.41 5.84 6.97 6.50
Cascaded ASR+GEC Methods
Gemma 2B (Team et al., 2024) FFT 13.20 12.30 10.40 1.60 2.60 2.20 3.70 6.00 7.50 6.61
Gemma 8x2B FFT 13.10 12.20 10.30 1.50 2.50 2.10 3.60 5.90 7.40 6.51
NeKo (Ours) Gemma 8x2B 13.00 12.10 10.20 1.40 2.40 2.00 3.50 5.80 7.30 6.41
Mistral 7B (Jiang et al., 2023) FFT 13.07 11.87 10.09 1.48 2.46 2.04 3.55 5.75 7.29 6.40
Mixtral 8x7B (Jiang et al., 2024b) FFT 12.91 12.19 10.34 1.54 2.55 2.12 3.64 5.89 7.43 6.51
Mixtral 8x7B Lora 12.96 12.24 10.38 1.55 2.56 2.13 3.66 5.92 7.47 6.60
Mistral 8x7B Branch-Train-Mix (Sukhbaatar et al., 2024) 13.13 11.93 10.14 1.49 2.47 2.05 3.57 5.78 7.33 6.43
NeKo (Ours) Mixtral 8x7B 12.55 11.82 10.02 1.49 2.47 2.05 3.52 5.76 7.25 6.34
NeKo (Ours) Mixtral 8x22B 12.61 11.93 10.15 1.52 2.51 2.09 3.58 5.82 7.33 6.40

### v2

# Yen-Ting Lin Zhehuai Chen Piotr Zelasko Zhen Wan Xuesong Yang Zih-Ching Chen Krishna C Puvvada Szu-Wei Fu  Ke Hu  Jun Wei Chiu  Jagadeesh Balam  Boris Ginsburg  Yu-Chiang Frank Wang  Chao-Han Huck Yang NVIDIA

@@ Abstract
Construction of a general-purpose post-recognition error corrector poses a crucial question: how can we most effectively train a model on a large mixture of domain datasets? The answer would lie in learning dataset-specific features and digesting their knowledge in a single model. Previous methods achieve this by having separate correction language models, resulting in a significant increase in parameters. In this work, we present Mixture-of-Experts as a solution, highlighting that MoEs are much more than a scalability tool. We propose a Multi-Task Correction MoE, where we train the experts to become an “expert” of speech-to-text, language-to-text and vision-to-text datasets by learning to route each dataset’s tokens to its mapped expert. Experiments on the Open ASR Leaderboard show that we explore a new state-of-the-art performance by achieving an average relative 5.05.0% WER reduction and substantial improvements in BLEU scores for speech and translation tasks. On zero-shot evaluation, NeKo outperforms GPT-3.5 and Claude-3.5 Sonnet with 15.515.5% to 27.627.6% relative WER reduction in the Hyporadise benchmark. NeKo performs competitively on grammar and post-OCR correction as a multi-task model.
[Uncaptioned image]NeKo: Cross-Modality Post-Recognition Error Correction with Tasks-Guided Mixture-of-Experts Language Model
Yen-Ting Linthanks: Work done at NVIDIA research as an intern.  Zhehuai Chen  Piotr Zelasko  Zhen Wan  Xuesong Yang Zih-Ching Chen  Krishna C Puvvada  Szu-Wei Fu   Ke Hu   Jun Wei Chiu   Jagadeesh Balam   Boris Ginsburg   Yu-Chiang Frank Wang   Chao-Han Huck Yang NVIDIA corresponding authors: ytl@ieee.org, hucky@nvidia.com
@@ 1 Introduction
Human recognition capabilities span multiple modalities, including speech recognition, visual patterns, and extensions to semantic and textual interpretations. These faculties, however, are not infallible and often incorporate mis-recognition errors. Despite these imperfections, humans efficiently communicate using speech, language, or facial expressions.
@@
For instance, two non-native speakers (Lev-Ari, 2015; Valaki et al., 2004) can often achieve mutual understanding through this imperfect recognition and subsequent interpretative processes, even when the conversation is marred by lexical inaccuracies and subdued accents. In other words, humans (as intelligent agents) exhibit a robust capacity for generative understanding Jiang et al. (2020); Cheng et al. (2021) that extends beyond initial recognition results. In neuroscience Zatorre and Gandour (2008), the inferior temporal gyrus and the temporal lobe are not confined to rudimentary perception but are also integral to the post-recognition processes that facilitate semantic understanding of language (Levinson and Evans, 2010), speech (Marshall et al., 2015), and visual patterns (Vink et al., 2020). This form of “post-recognition correction,” exemplified by the application of language modeling (LM) to initial recognition outputs, has been introduced to the field for both acoustic (automatic speech recognition, ASR) and visual (optical character recognition, OCR) modalities.
@@
With the LMs scaling up to LLMs (Brown et al., 2020), recent efforts (Chan et al., 2023; Yang et al., 2023; CHEN et al., 2023; Hu et al., 2024a) have focused on exploring a “generative modeling” for post-recognition correction. This generative error correction (GER) approach uses LLMs to conduct final recognition from given first-pass text-based predictions from recognition models, including ASR, image captioning (IC), and machine translation (MT). This cascaded two-agents text-to-text GER model has outperformed larger single multi-modal and multi-task models in these tasks. Meanwhile, these GER solutions heavily depends on domain-specific fine-tuning processes (Chen et al., 2024a) that utilize parameter-efficient components, which often suffers a performance degradation from a lack of generalizability across different datasets, domains, and tasks.
@@
To characterize “model generalization,” mixture-of-experts (MoE) (Jiang et al., 2024a) has emerged as a promising approach for multi-task learning, consisting of of a set of expert networks and a gating network that learns to route the input to the most appropriate expert (Sukhbaatar et al., 2024). This enables MoE models to learn more specialized and fine-grained representations compared to monolithic models. However, most MoE models are designed for general-purpose language modeling(Dai et al., 2024), with experts not explicitly assigned to specific tasks, but rather learn to specialize in different aspects of the input space through data-driven training. Effectively leverage MoE for multi-task error correction, where the experts need to capture task-specific features while allowing knowledge sharing, remains an open question.
@@
In this work, we propose NeKo, a “geNerative multi-tasK error correction” approach that leverages a pre-trained MoE model to drive diverse tasks and cross-domain knowledge, as shown in Figure 1. The key idea is to continuously pre-train MoE model on a mixture of error correction datasets, with each expert specializing in a specific domain. This task-guided MoE fine-tuning approach enables the experts to capture task-specific features while allowing knowledge sharing through the router. We further pursue this direction by modeling MoE on error correction and highlight the effectiveness and robustness of MoEs in learning from a mixture of correction datasets.
@@
NeKo captures the nuances of each task, benefiting from shared knowledge across experts. Evaluated on tasks such as ASR, ST, OCR, and unseen textual error correction (TEC), NeKo consistently outperforms baseline models, including Claude-3.5 Sonnet and GPT-3.5. It achieves state-of-the-art WER reduction on the Hyporadise benchmark and large-scale Open ASR Leaderboard (Srivastav et al., 2023). NeKo also significant improves in OCR error correction. Further analysis confirms its robust multi-task capabilities. In summary, the main contributions of this work include:
@@
    We introduce NeKo, a multi-task error correction LLM that leverages task-guided mixture-of-experts for diverse post-recognition correction tasks. To the best of our knowledge, this is the first work that explores the use of MoE for multi-task error correction.
@@
    NeKo has been studied under a new form of cross-modalities post-recognition correction evaluation, serving as strong open-source ASR, ST, OCR, and TEC baselines. Our results show that NeKo discovers new state-of-the-art performance in ASR as a multi-task correction model.
@@
    The NeKo models, newly created source datasets, and training processes are scheduled to open source under the CC BY-SA 4.0 license to support reproducibility in future research.
@@ 2.1 Mixture-of-Experts (MoE)
2.2 Tasks-Guided Auxiliary Expert Assignment
@@
where $\text{top1}=\operatorname*{arg\,max}_{j\neq f(T_{i})}G(x)_{j}$ is the index of the top-1 expert selected by the gating network, excluding the task-specific expert $f(T_{i})$.
@@ 2.3 Training Objective
where $x$ is the input sequence (e.g., ASR hypotheses, OCR output), $y$ is the target sequence (e.g., ground-truth transcription, corrected text), and $p(y|x,T_{i})$ is the probability of the target sequence given the input sequence and the task prompt (Figure 3.) By jointly training on multiple error correction datasets with task-guided expert assignment, NeKo learns to capture task-specific features while allowing for knowledge sharing across tasks through the shared gating network and other model components.
@@ 3.1 Training and Evaluation Datasets
For the translation error correction task, we use the subset of the HypoTranslate dataset (Hu et al., 2024b) for training and evaluation. This dataset includes translation from FLEURS (Conneau et al., 2022), CoVoST-2 (Wang et al., 2020), and MuST-C (Di Gangi et al., 2019), covering a range of languages such as Spanish, French, Italian, Japanese, Portuguese, Chinese, and Persian.
@@
For the optical character recognition (OCR) error correction task, we use the en-us portion of the OCR dataset (PleIAs, 2023), which contains newspaper texts from Chronicling America.
@@
For the textual error correction (TEC) task, we use a subset of the CoEdIT dataset (Raheja et al., 2023) from Grammarly, which contains 82K task-specific instructions for text editing.
@@ 3.2 Task-Specific Recognition Systems and Baselines
We compare against state-of-the-art ASR models, Whisper-V2-Large (Radford et al., 2022), Canary (NVIDIA, 2024) without applying GEC method. End-to-end ASR-LLM, SALM (Chen et al., 2024b), Qwen2-audio, and Gemini-2-Flash have been also compared. For all Cascaded ASR+GEC Methods, the task-specific system is the Canary model. This model transcribes the speech data and generate 5-best hypotheses for each utterance using temperature-based sampling (Ackley et al., 1985) with $p=0.3$. This allows us to capture a diverse set of potential transcriptions for each utterance, which can be fed into our error correction model.
@@
We compare our proposed method against two baselines: (1) the input text without any correction (denoted as Baseline) and (2) a Mixtral 8x7B model fine-tuned only on the respective dataset for each task (denoted as Mistral 8x7B Direct Finetune). This allows us to assess the effectiveness of our task-guided expert assignment approach in handling OCR and TEC errors, as its ability to leverage knowledge from multiple tasks to improve performance on individual tasks compared to direct fine-tuning on a single dataset.
Table 1: Cross-domain ASR correction results in zero-shot and few-shot settings on the Hyporadise benchmark (CHEN et al., 2023). We compare NeKo against GPT-4 Turbo and Claude-3.5 Sonnet in 0- and 5-shot settings. The baseline represents the WER of task-specific model Whisper-Large. The oracle results used in CHEN et al. (2023) (N-best and Compositional) provide an upper bound for the correction performance.
Domain Test Set Baseline GPT-3.5 Turbo Claude-3.5 Sonnet 0-shot w/ NeKo Oracle
@@
Specific
Scenario
WSJ-dev93 9.0 $8.5_{{\color[rgb]{0,.5,.5}-5.6\%}}$ $7.7_{{\color[rgb]{0,.5,.5}-14.4\%}}$ $8.2_{{\color[rgb]{0,.5,.5}-8.9\%}}$ $7.4_{{\color[rgb]{0,.5,.5}-17.8\%}}$ $8.6_{{\color[rgb]{0,.5,.5}-4.4\%}}$ $7.5_{{\color[rgb]{0,.5,.5}-16.7\%}}$ $\textbf{6.8}_{{\color[rgb]{0,.5,.5}-24.4\%}}$ 6.5 5.3
@@
Common
Noise
CHiME4-bus 18.8 $17.6_{{\color[rgb]{0,.5,.5}-6.4\%}}$ $16.2_{{\color[rgb]{0,.5,.5}-13.8\%}}$ $17.1_{{\color[rgb]{0,.5,.5}-9.0\%}}$ $15.7_{{\color[rgb]{0,.5,.5}-16.5\%}}$ $17.7_{{\color[rgb]{0,.5,.5}-5.9\%}}$ $15.9_{{\color[rgb]{0,.5,.5}-15.4\%}}$ $\textbf{14.5}_{{\color[rgb]{0,.5,.5}-22.9\%}}$ 16.8 10.7
@@
Speaker
Accent
MCV-af 25.3 $24.9_{{\color[rgb]{0,.5,.5}-1.6\%}}$ $23.6_{{\color[rgb]{0,.5,.5}-6.7\%}}$ $24.4_{{\color[rgb]{0,.5,.5}-3.6\%}}$ $23.0_{{\color[rgb]{0,.5,.5}-9.1\%}}$ $25.0_{{\color[rgb]{0,.5,.5}-1.2\%}}$ $23.3_{{\color[rgb]{0,.5,.5}-7.9\%}}$ $\textbf{21.0}_{{\color[rgb]{0,.5,.5}-17.0\%}}$ 23.6 21.7
@@
[…]
Table 2: ASR correction results on the Open ASR Leaderboard. We report the Word Error Rate (WER) for each dataset and the average across all 9 datasets. NeKo establishes a new state-of-the-art performance on the leaderboard, outperforming both end-to-end ASR methods and cascaded ASR+GEC approaches. We report the actual tuning parameter in parentheses (.) and the sum of the frozen Whisper results in front.
Model Inference Para. Avg. $\downarrow$ AMI Earnings22 Gigaspeech LS Clean LS Other SPGI Tedlium Voxp. MCV9
ASR or SpeechLMs: End-to-end Voice Understanding Models
Distil-Whisper-V2-L (Gandhi et al., 2023) 0.75B 8.31 14.65 12.12 10.31 2.95 6.39 3.28 4.30 8.22 12.60
Whisper-V2-L (Radford et al., 2022) 1.5B 8.06 16.82 12.02 10.57 2.56 5.16 3.77 4.01 7.50 10.11
Canary (NVIDIA, 2024) 2B 6.67 14.00 12.25 10.19 1.49 2.49 2.06 3.58 5.81 7.75
Bestow Speech LM (Chen et al., 2024c) 1.8B 6.50 12.58 12.86 10.06 1.64 3.07 2.11 3.41 5.84 6.97
Qwen2-Audio (Chu et al., 2024) 8B 7.43 - - - 1.6 3.6 - - - -
Gemini-2.0-Flash - 8.56 - - - - - - - - -
ASR+LLM: Frozen Whisper-v2-L (1.5B) + Voice Correction LMs
+ Gemma 2B (Team et al., 2024) FFT 3.5B (2B) 6.61 13.20 12.30 10.40 1.60 2.60 2.20 3.70 6.00 7.50
+ Gemma 8x2B FFT 3.5B (2B) 6.51 13.10 12.20 10.30 1.50 2.50 2.10 3.60 5.90 7.40
NeKo (Ours) Gemma 8x2B 3.5B (2B) 6.41 13.00 12.10 10.20 1.40 2.40 2.00 3.50 5.80 7.30
NeKo (Ours) Qwen1.5-MoE 4.2B (2.7B) 5.90 12.60 11.82 9.95 1.30 2.32 1.94 3.20 5.80 7.30
+ Mistral 7B (Jiang et al., 2023) FFT 8.5B (7B) 6.40 13.07 11.87 10.09 1.48 2.46 2.04 3.55 5.75 7.29
+ Mixtral 8x7B (Jiang et al., 2024b) FFT 8.5B (7B) 6.51 12.91 12.19 10.34 1.54 2.55 2.12 3.64 5.89 7.43
+ Mixtral 8x7B Lora 8.5B (7B) 6.60 12.96 12.24 10.38 1.55 2.56 2.13 3.66 5.92 7.47
+ Mistral 8x7B BTM (Sukhbaatar et al., 2024) 8.5B (7B) 6.43 13.13 11.93 10.14 1.49 2.47 2.05 3.57 5.78 7.33
NeKo (Ours) Mixtral 8x7B 8.5B (7B) 6.34 12.55 11.82 10.02 1.49 2.47 2.05 3.52 5.76 7.25
NeKo (Ours) Mixtral 8x22B 23.5B (22B) 6.40 12.61 11.93 10.15 1.52 2.51 2.09 3.58 5.82 7.33